UT Apr 29 UT 2021 - Engineering Night



Request Forms for Night Use


Engineering Tasks Submitted

  • TCS Reflective Memory Test
    • recording the tcs reflective memory. It works independently from the TCS
    • run it with hopes of capturing information for IT 7380 or 7218
  • MODS
    • Primary Spectrophotomteric Standards - 1 hour
      • Any seeing, clear (photometric)
      • Cycle through all modes with each MODS
    • MODS Field Correction check for IT 8315 - 1-2hrs depending on number of targets
      • Seeing < 1.2"
      • Up to 1-2 mags of extinction
      • run IDL task field_collect (same as measuring field aberrations: ComTools#Field_Data_Collection)
      • do LDG and RDG in parallel (two IDL windows)
      • Use 6-8 stars (some fields only have 6 stars; start the list a second time, if you are under an hour)
        • Stone_G is at RA=10:17
        • Stone_H at RA=11:03
        • Stone_J at RA=13:32,
        • and other Stone / BS fields are available if earlier or later.
  • LUCI-AO - 1hour
    • update AOS software
      • close loops on several targets to check software update
      • AO or ESM
    • ESM/AO Crossover test
  • LBTI SX Com tasks - 0.5 night
    • First half of night

Night Plan

  • Activate TCS Reflective Memory recorder
  • LBTI Engineering Tasks w/Arcetri until 00:12
  • LUCI-1 AO Testing - until 03hr
    • Activate AO Software after handover
    • OSCO - until 01hr
      • this will test Bin 1
    • Targets for Bin 2 and Bin 3 to verify AO software update - 0.5 hr
      • Bin 2 Target
        • Candidate: On-axis NS target Davida - awaiting confirmation (CV on time for observations)
      • Bin 3 Target
        • Candidate: Mrk 231 - core is on-axis AO Ref Star (r~13.7) - but H/K are 10.4/9.2 mag - might need either H2/BrGamma or sub-frame
        • has been done with PUEO (FWHM~0.11") but not recently - HST/NIC1 only other obserations - Arcetri predicts DL FWHM~0.06"
    • AO/ESM Crossover test
      • This can also double for testing Bins - higher priority than Bin 2 and Bin 3 target above
      • Observe using the N375 camera but with full AO.
      • Stellar field with AO Ref star roughly in the center of the field
      • Repeat observation with ESM to compare
      • Goal: Determine the point at which AO corrections degrade significantly and are poorer than ESM correections
        • M92 R=9.4 (WFSmag=10.7 - Bin 1)
        • M92 R=15.1 (WFSmag forced to 16.51 to test Bin 4)
    • ESM On-Axis AO Ref Star test
      • Hoag's Object - On-axis AO Ref star (using core of the galaxy) - 1.5hr
        • RA 15hr Dec +21
        • WFSmag estimate = 14.7
        • elevation 45 at twilight
        • this will test Bin 4
  • If time - MODS Field Correction Check
    • Stone L or Stone M best in RA

Contingency

  • If first half lost to weather - continue 1-2 hours for LBTI then reconfigure for LUCI-1 AO

Poor Weather Backups - 2nd half

  • Velocity Dispersion of local mergers - LD-2019B-005 - 1hr per target bino-MODS
    • NGC 6598 - all up 2nd half of night
    • any conditions - go longer if poorer conditions
  • Velocity Dispersions of local mergers - LD-2019B-005 - 1hr per target bino-LUCI
    • Any conditions - go longer if poorer conditions
    • NGC 6598 - all up 2nd half of night
    • NGC 4004 - sets 3am
    • UGC 10607
    • UGC 10675

Summary

AO log found here: https://wiki.lbto.org/AdaptiveOptics/20210429_SX_Night

0:30 Fog on the summit. No snow or rime ice build up on the dome. If humidity and clouds give way david is confident we can open.

1:30 Amali and Steve are online and set up. The clouds are giving way but humidity is pegged. Winds are ~13m/s

3:15 Humidity has dropped to 95%. Some clouds overhead. Winds up to 16m/s. I've init the lucis and parked the blind masks for later btw. Used my fancy new fieldstop check script to adjust the stop a few pixels on luci1

4:54 We have the trifecta: humidity pegged, clouds, and now winds over the limits.

07:12 - nominal handover time - not open - still pegged at humidity limit, winds just at limit - clouds thick and fog

07:16 - utility chiller in alarm last few hours - mirror ventilation stopped a few hours ago. DG.Huerta noted this went into alarm around UT 03hr (8pm local time).

08:31 - B.Bingham was able to power cycle but drives up the return temperature.

08:35 - clouds thinning slightly (sucker hole overhead - but humidity remains at 99% with winds close to the limits in the last 15min.

09:01 - utility chiller issue still a problem. mirror ventialtion is down to reduce the load on it.

-- BarryRothberg - 07 Apr 2021
Topic revision: r15 - 29 Apr 2021, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback